MEASUREMENTS OF PEROXY RADICALS USING CHEMICAL AMPLIFICATION/CAVITY RING-DOWN SPECTROSCOPY

YINGDI LIU, Department of Chemistry, University of California, Riverside, CA 92521; RODRIGO MORALES-CUETO, Instituto de Quimica, Universidad Nacional Autonoma de Mexico, Ciudad Universitaria, 04510, Mexico, D.F., Mexico; JAMES HARGROVE, DAVID MEDINA, Department of Chemistry, University of California, Riverside, CA 92521; and \underline{JINGSONG ZHANG}, Department of Chemistry and Air Pollution Research Center, University of California, Riverside, CA 92521.

The peroxy radical chemical amplification (PERCA) method is combined with cavity ring-down spectroscopy (CRDS) to measure peroxy radicals HO_{2} and RO_{2}. HO_{2} and RO_{2} are converted to NO_{2} via reactions with NO, with subsequent reactions to recycle most of the OH and RO coproducts back to HO_{2} by reactions with CO and O_{2} and amplify the level of NO_{2}. The amplified NO_{2} is then monitored by CRDS, a sensitive absorption technique. The PERCA-CRDS method is calibrated with a HO_{2} radical source (0.5-3 ppb); using a 3-meter Teflon tubing as the reactor and 2.5 ppmv NO and 5% vol/vol CO, the PERCA amplification factor or chain length is determined to be 150 +-50 in this study. The peroxy radical detection sensitivity by PERCA-CRDS is approx. 5 pptv/30s. Ambient measurements of the peroxy radicals are carried out at Riverside, California in 2007 to demonstrate the PERCA-CRDS method.